OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 1 + 3i A. 1 + 9i B. 10 C. -8 D. 1 - 9i

OpenStudy (anonymous):

i thought it was D. but i just want to see if im right or not?

OpenStudy (kropot72):

It is not D. Did you multiply it out using the FOIL method? (1 + 3i)(1 - 3i) = ?

OpenStudy (anonymous):

oh no i didnt.. hmm ..

OpenStudy (anonymous):

\[(a+bi)(a-bi)=a^2+b^2\] a real number don't think about \(i\) and don't think about minus signs, just straight up \(a^2+b^2\)

OpenStudy (anonymous):

would it be 10?

OpenStudy (anonymous):

and i might add no foil

OpenStudy (kropot72):

Yes , the answer is 10.

OpenStudy (anonymous):

yes, \(1^2+3^2=1+9=10\)

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

would either of you know how to do this problem; Find the real numbers x and y that make the equation true. -3 + yi = x + 6i

OpenStudy (kropot72):

Hint: If y = 6 then the imaginary terms cancel.

OpenStudy (anonymous):

so would y= 6 x= -3

OpenStudy (calculusfunctions):

Sure! In any equation, left side equals to the right side iff each component on one side of the equation equals to its corresponding component on the other side.

OpenStudy (kropot72):

@katherinekc Yes, you are right.

OpenStudy (calculusfunctions):

Therefore if you have two complex numbers equal to each other then the real part of the complex number on one side equals to the real part of the other complex number on the other side of the equation, and the imaginary part of the complex number on one side equals to the imaginary part of the complex number on the other side of the equation.

OpenStudy (anonymous):

so then id be correct when saying y= 6 x= -3

OpenStudy (calculusfunctions):

The real parts on both sides are -3 and x. Therefore x = -3. The imaginary parts on both sides are yi and 6i. Therefore y = 6. Therefore you are one hundred percent correct @katherinekc! Great work!

OpenStudy (anonymous):

thankyou!!

OpenStudy (calculusfunctions):

Of Course! You deserve it!